# Knowledge Graph Large Language Model (KG-LLM) for Link Prediction

### 13 Mar 2024 v2

@@ 5.4 Dataset 1
5.4.1 Standard Framework One-shot In-Context Learning for Multi-hop Link Prediction
@@ 5.4.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
5.4.3 Standard Framework One-shot In-Context Learning for Multi-hop Relation Prediction
@@ 5.5 Dataset 2
5.5.1 Standard Framework One-shot In-Context Learning for Multi-hop Link Prediction
@@ 5.5.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
5.5.3 Standard Framework One-shot In-Context Learning for Multi-hop Relation Prediction
@@ 5.6 Examples
Answer from Standard FLAN-T5-Large Model
@@
Answer from Standard FLAN-T5-Large Model
@@
Answer from Standard FLAN-T5-Large Model
@@
Answer from Standard FLAN-T5-Large Model
@@
Table 6: Standard FLAN-T5-Large Model Responses to Modified Inputs
@@ 1. Yes
Answer from Standard LlaMa2-7B Model
@@
Answer from Standard LlaMa2-7B Model
@@
Answer from Standard LlaMa2-7B Model
@@
Answer from Standard LlaMa2-7B Model
@@
Table 8: Standard LlaMa2-7B Model Responses to Modified Inputs
@@
Answer from Standard Gemma-7B Model
@@
Answer from Standard Gemma-7B Model
@@
Answer from Standard Gemma-7B Model
@@
Answer from Standard Gemma-7B Model
@@
Table 10: Standard Gemma-7B Model Responses to Modified Inputs

### 18 Mar 2024 v3

@@ 5.4 Dataset 1
5.4.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
@@ 5.4.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
5.4.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
@@ 5.5 Dataset 2
5.5.1 ablation Framework One-shot In-Context Learning for Multi-hop Link Prediction
@@ 5.5.2 KG-LLM Framework One-shot In-Context Learning for Multi-hop Link Prediction
5.5.3 ablation Framework One-shot In-Context Learning for Multi-hop Relation Prediction
@@ 5.6 Examples
Answer from ablation FLAN-T5-Large Model
@@
Answer from ablation FLAN-T5-Large Model
@@
Answer from ablation FLAN-T5-Large Model
@@
Answer from ablation FLAN-T5-Large Model
@@
Table 6: ablation FLAN-T5-Large Model Responses to Modified Inputs
@@ 1. Yes
Answer from ablation LlaMa2-7B Model
@@
Answer from ablation LlaMa2-7B Model
@@
Answer from ablation LlaMa2-7B Model
@@
Answer from Ablation LlaMa2-7B Model
@@
Table 8: Ablation LlaMa2-7B Model Responses to Modified Inputs
@@
Answer from Ablation Gemma-7B Model
@@
Answer from Ablation Gemma-7B Model
@@
Answer from Ablation Gemma-7B Model
@@
Answer from Ablation Gemma-7B Model
@@
Table 10: Ablation Gemma-7B Model Responses to Modified Inputs